how gravitational acceleration varies with height .
1
Expert's answer
2020-09-15T10:05:03-0400

Gravitational acceleration depends on height as

"\displaystyle g_h = g_0 (\frac{R_e}{R_e+h})^2" ,

where "g_h" is gravitational acceleration at a certain height, "g_0 = 9.80665"  m/s2 is standard gravitational acceleration, "R_e" is the Earth's radius, "h" is height.

As we see from the formula, gravitational acceleration decreases with height.

So when "h=0" m, that means that you're at the Earth ocean-level surface, then "g_h = g_0". In mountains "g_h" will be just a bit smaller, the difference is less than 1 %. Only going to the height of ionosphere and low orbit satellites ("h \leq 2000" km) will make you feel a big difference.